OpenStudy (anonymous):

MEDAL!! The following statement describes the number of marbles, n, in a bag. Ten less than eight times the number of marbles is greater than or equal to the sum of six times the number of marbles and fourteen. Create an inequality to represent this statement, and solve for n. Then, complete the following sentences about the solution. To graph the solution of the inequality on a horizontal number line, first place (Look at #1) . Then, shade the number line to the (Look at #2 . The solution inequality represents that the number of marbles in the bag is (Look at #3)(Look at #4)

OpenStudy (anonymous):

#1 A. Open Circle at 12 B. Closed Circle at 12 C. Open circle at 24 D. Closed circle at 24

OpenStudy (anonymous):

#2. A. Right B. Left

OpenStudy (anonymous):

#3 A. Less Than B. More Than C. No less than D. No More Than

OpenStudy (anonymous):

#4 A. 10 B. 12 C. 14 D. 24

OpenStudy (anonymous):

@iGreen @texaschic101

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (texaschic101):

8n - 10 >= 6n + 14 8n - 6n >= 14 + 10 2n >= 24 n >= 12 open circle on 12 because there is an equal sign...greater means shaded to the right...more then (or equal to)12

OpenStudy (anonymous):

So what is the answer? Is it B, A, D, D

OpenStudy (texaschic101):

wait..I messed up.....first one is closed circle because it has an equal sign....so it goes... A,A,B,B

OpenStudy (anonymous):

It was wrong, but no worrys. It wasn't for a grade. :) Thanks anyway! I will medal you for helping me get at least half of them right. :)